Why is f not a function from R to R if

a) f (x) = 1/x?

b) f (x) =√x?

c) f (x) = ±√(x^2+1)?


1
Expert's answer
2021-09-02T00:38:00-0400

A function "f" from "\R" to "\R" is a rule that assigns to each element "x\in \R" exactly one

element, called "f(x), f(x)\in \R." in

a) "f(x)=\dfrac{1}{x}"

"x\not=0"

The function "f" is undefined at "x=0."

Therefore a function "f(x)=\dfrac{1}{x}" is not a function from "\R" to "\R."


b) "f(x)=\sqrt{x}"

"x\geq0"

The function "f" is undefined for "x\in \R,x<0."

Therefore a function "f(x)=\sqrt{x}" is not a function from "\R" to "\R."


c) "f(x)=\pm\sqrt{x^2+1}"

"f(0)=\pm\sqrt{0^2+1}=\pm1"

A relation "f(x)=\pm\sqrt{x^2+1}" assigns to "x=0" two elements "-1" and "1."


Therefore "f(x)=\pm\sqrt{x^2+1}" is not a function from "\R" to "\R."